OpenStudy (anonymous):

Solve the equation. 4 over x^2 − 2x − 35 = 5 over x^2 + 9x + 20 x=

OpenStudy (anonymous):

do you only want an answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

PLEASE

OpenStudy (anonymous):

The roots will be imaginary

OpenStudy (anonymous):

KK

OpenStudy (anonymous):

-11 + SQRT(99) i /2 ,-11 - SQRT(99) i /2

OpenStudy (anonymous):

that's the answer?

OpenStudy (anonymous):

They are the roots of euqtion

OpenStudy (anonymous):

x=51 ?

OpenStudy (anonymous):

x^2 +11x +55 = 0 D = 11^2 - (4*55) = 121 - 220 which is negative hence roots are imaginary

OpenStudy (anonymous):

This is a place to learn, not get free answers.

OpenStudy (anonymous):

Is the eq: (4)/(x^2-2x-35) = (5)/(x^2+9x+20) If so, you could be able to factor both eq... elminate (x-5) and solve for x, in this case, x=51.

OpenStudy (anonymous):

lol. im writing all their notes.....@LAZY

OpenStudy (anonymous):

Sorry >.>

OpenStudy (anonymous):

Move the RHS to the LHS.\[\frac{4}{x^2-2 x-35}-\frac{5}{x^2+9 x+20}=0 \]Then combine the fractions.\[\frac{51-x}{(x-7) (x+4) (x+5)}=0 \]By inspection, x = 51 .

OpenStudy (anonymous):

KOOL:) THANK YOU!

OpenStudy (anonymous):

Thank you for the medal.